what volume of 0.1M KMnO4 is need to oxidise 100mg of FeC2O4 in acidic medium?

Grade:11

2 Answers

RON
9 years ago
As their number of equivalents are equal i.e n1v1=n2v2
so,
We just need to find out these variables.
n1 i.e normality of KMn04= molarity x n-factor
n-factor of KMn04 in acidic medium is 5
normality = 0.1(given molarity) x 5
=0.5
so n1v1= 0.5x v1
Now to find out equivalents of Fec204 we have to to use the formula n.eq=given mass/eq.mass.
Given mass=.1g
eq.mass=mol. mass/n-factor
mol mass= roughly 144
n-factor of FeC204 in acidic medium=1
therefore. eq. mass=144
.Hence number of gram equivalents=0.1/144.
Now according to the law of equivalence .
n1v1=n2v2 (equal gram equivalents)
0.5xv1=0.1/144
v1=1/(144*5)
=1/720.
therefore our v1 comes out to be 0.00139 litres or 1.39 ml.
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

As their number of equivalents are equal i.e n1v1=n2v2
so,
We just need to find out these variables.
n1 i.e normality of KMn04= molarity x n-factor
n-factor of KMn04 in acidic medium is 5
normality = 0.1(given molarity) x 5
=0.5
so n1v1= 0.5x v1
Now to find out equivalents of Fec204 we have to to use the formula n.eq=given mass/eq.mass.
Given mass=.1g
eq.mass=mol. mass/n-factor
mol mass= roughly 144
n-factor of FeC204 in acidic medium=1
therefore. eq. mass=144
.Hence number of gram equivalents=0.1/144.
Now according to the law of equivalence .
n1v1=n2v2 (equal gram equivalents)
0.5xv1=0.1/144
v1=1/(144*5)
=1/720.
therefore our v1 comes out to be 0.00139 litres or 1.39 ml.

Thanks and Regards